# Collective dynamics versus entanglement in quantum battery performance

### v2

@@ Abstract
Identifying the physical origin of enhanced charging performance in many-body quantum batteries is a key challenge in quantum thermodynamics. We investigate whether improvements in stored energy and instantaneous charging power arise from genuine quantum correlations or from coherent collective dynamics that are not intrinsically quantum. We compare the time evolution of energetic quantities with a hierarchy of information-theoretic measures probing bipartite, tripartite, and further-partite correlations. Across different battery–charger configurations, we find a consistent temporal ordering in which the instantaneous power peaks before the buildup of strong quantum correlations, indicating that peak charging is dominated by coherent transport, while entanglement and scrambling develop at later times. Furthermore, charging protocols based on $\kappa$-local interactions are examined under both unconstrained and norm-constrained (fair) settings, enabling a clear distinction between classical scaling effects and genuine collective enhancements. Increasing the interaction order or the participation number does not automatically translate into higher charging power. Instead, the performance is primarily dictated by how many particles actually become mutually correlated and contribute to entanglement. Fully collective interactions provide a genuine advantage because all particles participate coherently, whereas partially extended interaction schemes fail to monotonically increase the number of effectively interacting particles, and therefore do not guarantee improved charging efficiency.
@@ I Introduction
Quantum batteries (QBs) are devices engineered to store and deliver energy through controlled quantum dynamics and have emerged as a central platform in quantum thermodynamics and quantum technologies. Importantly, QBs can be realized using quantum many-body systems, such as interacting spin chains, where collective effects, many-body interactions, and correlations can enhance charging power and lead to nontrivial scaling behaviors [alicki2013entanglement, binder2015quantacell, campaioli2017enhancing, ferraro2018high, andolina2019quantum, andolina2019extractable, le2018spin, rodriguez2021collective]. Beyond conventional spin-chain architectures, QBs have been investigated in topological systems as well as in various spin-based models subject to different types of external driving or interactions, including Floquet-engineered setups, kicked-spin models, and long-range interacting frameworks. These studies highlight the roles of periodic driving, operator spreading, and topological protection in the storage and transfer of energy [Puri2024FloquetManyBody, lu2025topological, Puri2024FloquetManyBody, Romero2025, sahoo2025power].
Moreover, QBs have been extensively investigated in open quantum system scenarios, where environmental effects such as dissipation, decoherence, and engineered reservoirs strongly influence charging efficiency, stability, and extractable work. These studies reveal phenomena including noise-assisted and dissipative charging, reservoir-mediated energy transfer, non-Markovian memory effects, thermal and composite environment influences, and fundamental power bounds [Ferraro2018, Andolina2019, GarciaPintos2020, Quach2020, Tabesh2020EnvMediated, Morrone2022NonMarkov, EnhanceOpenQB2023, Yao2025Reservoir, QuantumErgotropy2025, NonMarkovNSpin2025, Xu2023CompositeEnv, Zhang2024OpenInteractive, Zhang2024OpenEnvPerformance, EntropicUncertaintyQB2022, liu2025open, mathew2025super]. Collectively, these works emphasize that both the internal structure of the quantum battery and the surrounding environment critically determine charging performance.
From a structural perspective, quantum battery architectures are commonly categorized into two paradigms: (i) interacting batteries charged by noninteracting systems [le2018spin, ghosh2020enhancement, zhao2021quantum], and (ii) noninteracting batteries charged by interacting many-body chargers [rossini2020quantum, ferraro2018high, julia2020bounds]. In both cases, charging performance depends sensitively on the Hamiltonian structure, interaction range, and normalization, with rigorous bounds linking charging power to the Hamiltonian norm [julia2020bounds].
A central goal in quantum battery research is to identify genuinely quantum mechanisms such as coherence, entanglement, and collective dynamics that can enhance stored energy, extractable work, and especially charging power beyond classical limits [gyhm2024beneficial]. Early studies demonstrated that global collective driving can enhance instantaneous charging power relative to parallel local protocols [alicki2013entanglement, binder2015quantacell, campaioli2017enhancing], motivating the idea that entanglement and multipartite correlations might serve as key resources for fast charging. However, subsequent investigations clarified that entanglement is neither necessary nor sufficient for charging enhancement [andolina2019quantum, andolina2019extractable, julia2020bounds, kamin2020entanglement, shi2022entanglement], as maximal charging power is often reached prior to the development of strong multipartite correlations [kamin2020entanglement, shi2022entanglement, perarnau2015extractable].
A key challenge in the study of many-body quantum batteries is identifying the physical origin of enhanced charging performance. In particular, it is crucial to understand whether improvements in stored energy and instantaneous charging power arise from genuinely quantum correlations or from coherent collective dynamics that are not intrinsically quantum. To address this question, we analyze the temporal evolution of energetic quantities alongside a hierarchy of information-theoretic measures probing bipartite, tripartite, and multipartite correlations. Across different battery–charger configurations, we find a consistent temporal ordering: the instantaneous power peaks before the buildup of strong quantum correlations, indicating that coherent transport dominates peak charging, while entanglement and scrambling develop at later times.
We further investigate charging protocols based on $\kappa$-local interactions under both unconstrained and norm-constrained (fair) conditions, which allows us to separate classical scaling effects from genuine collective behavior. Our analysis shows that increasing the interaction order or participation number alone does not guarantee enhanced charging power. Instead, the performance of the battery is governed by the coherence and structure of the interaction network: fully collective interactions provide a genuine advantage, whereas partially extended interaction schemes can suppress power due to competing correlation pathways. These findings clarify the distinct roles of coherence, entanglement, and interaction structure in optimizing the charging performance of many-body quantum batteries.
@@
The paper is organized as follows. Section II introduces the quantum battery and charger Hamiltonians and defines measures for bipartite, tripartite, and multipartite entanglement. Section III presents the charging dynamics for both configurations—interacting battery with noninteracting charger, and noninteracting battery with interacting charger—and examines the effects of $\kappa$-local interactions and interaction range on energy transfer and correlations. Section IV summarizes the main findings and concludes the study.
@@
Quantum Battery and Charger Hamiltonian
@@
To initiate the charging process, we switch on the interactions among the battery spins at time $t=0$. Additionally, we include an external field applied in the $x$-direction, which further drives the charging dynamics. The resulting evolution is governed by the interacting spin Hamiltonian
@@
where $J_{\alpha}$ are the nearest-neighbor interaction strengths in the $\alpha$-direction, and $h_{x}$ is the magnetic field applied to each spin in the x-direction. $\kappa$ quantifies the number of interacting spins, representing the degree of $\kappa$-locality. By varying $\kappa$, we examine how higher-order spin correlations influence charging performance and the generation of multipartite entanglement. This Hamiltonian simultaneously induces spin correlations and drives the energy injection that charges the battery. The charging dynamics of the battery are governed by the Hamiltonian
@@
Concurrence is one of the most widely used measures of bipartite entanglement for two-qubit systems [wootters1998entanglement]. It provides a quantitative value ranging from 0 for separable (uncorrelated) states to $1$ for maximally entangled states. In the context of quantum batteries, concurrence is useful for quantifying the entanglement between specific pairs of spins, for example, the first and last spins of a chain, which can reveal the formation of long-range correlations during the charging process.
For a two-qubit density matrix $\rho$, we define the spin-flipped matrix as $\tilde{\rho}=(\sigma_{y}\otimes\sigma_{y})\rho^{*}(\sigma_{y}\otimes\sigma_{y})$, where $\rho^{*}$ is the complex conjugate in the computational basis and $\sigma_{y}$ is the Pauli-$y$ matrix. Using this, we construct $R=\rho\tilde{\rho}$. Let $\{\mu_{i}\}_{i=1}^{4}$ be the eigenvalues of $R$ (which are always real and nonnegative), and let their square roots be arranged in decreasing order: $\lambda_{1}\geq\lambda_{2}\geq\lambda_{3}\geq\lambda_{4}\geq 0$. The concurrence is then given by
$C(\rho)=\max\Big(0,\lambda_{1}-\lambda_{2}-\lambda_{3}-\lambda_{4}\Big).$
@@
Bipartite entanglement entropy (BEE) quantifies the degree of entanglement between a subsystem $X$ and its complement $X^{c}$ in a many-body system [Calabrese2005, Page1993]. It is computed from the entropy of the reduced density matrix associated with the subsystem,
@@
To study correlations among three subsystems simultaneously, we use the tripartite mutual information (TMI), defined as [iyoda2018scrambling, seshadri2018tripartite, Shukla2025ScramblingIsing]
@@
[…]
The QFI quantifies how sensitive the battery state is to unitary transformations generated by $\hat{H}_{B}$, and larger QFI indicates stronger collective correlations across spins, which can enhance energy storage and transfer in quantum batteries.
Furthermore, QFI can witness multipartite entanglement. For an $N$-spin battery, if the inequality
$F_{Q}[\hat{H}_{B}]\leq\Big\lfloor\frac{N}{\kappa}\Big\rfloor\kappa^{2}+\Big(N-\Big\lfloor\frac{N}{\kappa}\Big\rfloor\kappa\Big)^{2}$ (9)
[…]
@@
While BEE measures correlations for a single partition, the average bipartite entanglement entropy (ABEE) provides a global view of entanglement across the system. It is computed by averaging the BEE over all contiguous bipartitions:
@@
Figure 1: Noninteracting battery charged by an interacting Hamiltonian. Bipartite entanglement $(a_{1},b_{1},c_{1})$: Time evolution of the instantaneous power $P_{i}$, concurrence between the first and last spins $C(1,N)$, and bipartite entanglement entropy (BEE) $S_{N/2}$. Tripartite entanglement $(a_{2},b_{2},c_{2})$: Time evolution of $P_{i}$ and tripartite mutual information (TMI) $I_{3}$. Multipartite entanglement $(a_{3},b_{3},c_{3})$: Time evolution of $P_{i}$, quantum Fisher information (QFI) $F_{Q}$, and average entanglement entropy (AEE) $\overline{S}_{N/2}$. Panels correspond to: $(a_{1},a_{2},a_{3})$ $J_{x}=1$, $J_{y}=J_{z}=0$; $(b_{1},b_{2},b_{3})$ $J_{x}=J_{y}=1$, $J_{z}=0$; $(c_{1},c_{2},c_{3})$ $J_{x}=J_{y}=J_{z}=1$. Number of spins: $N=8$ and $h_{x}=h_{z}=1$.
@@ III Results
Noninteracting Battery and Interacting Charger
We first consider a configuration in which the battery Hamiltonian, $H_{B}$, represents a noninteracting system, and the same physical system also serves as the charger through the Hamiltonian $H_{C}$. The interaction terms are switched on only during the charging interval $0,\tau]$, at which point the charger becomes active and drives the dynamics. Its internal couplings mediate the energy flow into the battery and simultaneously generate quantum correlations. In this framework, the noninteracting battery acts as a passive energy receptor, allowing us to clearly isolate and analyze the role of charger-induced correlations in the charging process.
Figure 2: Interacting battery charged by an noninteracting Hamiltonian. Bipartite entanglement $(a_{1},b_{1},c_{1})$: Time evolution of the instantaneous power $P_{i}$, concurrence between the first and last spins $C(1,N)$, and bipartite entanglement entropy (BEE) $S_{N/2}$. Tripartite entanglement $(a_{2},b_{2},c_{2})$: Time evolution of $P_{i}$ and tripartite mutual information (TMI) $I_{3}$. Multipartite entanglement $(a_{3},b_{3},c_{3})$: Time evolution of $P_{i}$, quantum Fisher information (QFI) $F_{Q}$, and average entanglement entropy (AEE) $\overline{S}_{N/2}$. Panels correspond to: $(a_{1},a_{2},a_{3})$ $J_{x}=1$, $J_{y}=J_{z}=0$; $(b_{1},b_{2},b_{3})$ $J_{x}=J_{y}=1$, $J_{z}=0$; $(c_{1},c_{2},c_{3})$ $J_{x}=J_{y}=J_{z}=1$. Number of spins: $N=8$ and $h_{x}=h_{z}=1$.
@@
We investigate the relationship between the battery charging rate and bipartite correlations. To quantify bipartite correlations, we evaluate two complementary measures: the concurrence and the BEE. These quantities are computed for three representative charger interaction configurations: ($a_{1}$) $J_{x}=1$, $J_{y}=0$, $J_{z}=0$; ($b_{1}$) $J_{x}=1$, $J_{y}=1$, $J_{z}=0$; and ($c_{1}$) $J_{x}=1$, $J_{y}=1$, $J_{z}=1$. In all cases, the field strength is fixed to $h_{x}=1$. Concurrence quantifies pairwise entanglement between two qubits; here, we focus on the first and last spins to probe long-range correlations. In contrast, the BEE measures the entanglement between a chosen subsystem and its complement through the BEE of the reduced density matrix. In our calculations, we consider equal bipartitions of the spin chain. For a meaningful comparison, all three quantities are normalized by their respective maximum values, allowing us to identify the time regimes in which each quantity attains its peak.
@@
During the initial charging stage the instantaneous power $P_{i}(t)$ rises rapidly as the charger injects excitations into the battery; this rise is driven by coherent excitation transfer enabled by the charger’s internal couplings. Concurrence and BEE also increase during this phase, indicating that entanglement participates in the energy transfer process. Crucially, however, the peak of instantaneous power is reached before concurrence and BEE attain their maxima (see Fig. 1 $(a_{1},b_{1},c_{1})$). Concurrence typically grows faster than BEE and peaks earlier than BEE, but both lag the power peak. Physically, this temporal ordering shows that short-time coherent transport, which governs peak power, precedes the formation of fully developed bipartite entanglement across the system. Entanglement is therefore a necessary resource for enhanced charging, but it is not the sole determinant of the timing of maximal power: the charger’s coupling strengths and the coherent transfer channels set the time scale for the power peak, whereas entanglement continues to build up afterwards as correlations spread.
@@
To probe correlations beyond pairwise contributions, we evaluate the TMI, which quantifies how information and correlations are distributed among three subsystems. For this calculation, the spin chain of length $N$ is partitioned into four segments of size $N/4$ in total, and the initial state is taken to be the ground state of the battery Hamiltonian. This ensures consistency between the injected energy and the resulting correlation production during the charging process. The TMI increases substantially only after the early-time rise of energy and power, and its maximum occurs at later times compared to energetic quantities [Fig. 1 $(a_{2},b_{2},c_{2})$)]. This delayed growth of TMI reveals that three-way and higher-order correlations require additional time to form: pairwise coherences must first be established and propagated before genuine three-body information sharing emerges. From a physical standpoint, this hierarchical buildup reflects the propagation and interference of excitations through the interacting charger, leading ultimately to delocalized correlation structures (scrambling) that are slower to develop than local energy transfer.
@@
Energy injection from an interacting charger initially excites coherent transport channels that rapidly transfer energy into the battery and produce a pronounced peak in instantaneous power. The formation of quantum correlations pairwise, tripartite, and finally multipartite follows this energetic surge because correlations require the accumulation of relative phases and interaction-mediated spreading of excitations. Consequently, entanglement and scrambling develop on longer timescales than energy transfer. Therefore, although entanglement is essential to realize quantum advantages in charging (for instance by enabling nonclassical interference pathways or collective enhancement), there is no one-to-one temporal correspondence between entanglement maxima and the peak of instantaneous power. Instead, optimal charging occurs in an intermediate regime where coherence and modest entanglement combine to maximize power, while more extensive entanglement emerges later and contributes to the global reorganization and stabilization of the stored energy.
Interacting Battery and Noninteracting Charger
We now consider the reverse configuration, where the battery is interacting and the charger is noninteracting, effectively interchanging the roles of the two subsystems studied earlier. This setup allows us to assess how internal interactions within the battery itself influence the charging process and the buildup of correlations. Similar to the previous case, we monitor the time evolution of the instantaneous power, stored energy, and various entanglement measures to understand how the interacting battery dynamics unfold.

### v3

@@ Abstract
Identifying the origin of enhanced charging performance in many-body quantum batteries remains a central challenge in quantum thermodynamics. It is unclear whether improvements in stored energy and instantaneous charging power stem from genuinely quantum correlations, such as entanglement, or from coherent collective dynamics, in which energy is transferred through the battery by many particles acting together in a coordinated, phase-preserving manner. Here, we address this question by comparing the time evolution of energy and a hierarchy of entanglement measures probing bipartite, tripartite, and multipartite correlations. Across diverse battery–charger configurations, the instantaneous power peaks early, before significant entanglement develops, indicating that peak charging is dominated by coherent collective transport. Further analysis of $\kappa$-local interactions under fair constraints shows that only fully collective schemes ($\kappa=N$) engage all particles, aligning entanglement growth with energy storage and yielding a genuine enhancement. Partially extended interactions ($\kappa<N$) leave many particles inactive and fail to improve performance. Our analysis indicates that the charging advantage arises not from entanglement alone, but from correlations that coherently involve the entire system.
@@ I Introduction
Quantum batteries (QBs) are devices designed to store and deliver energy through controlled quantum dynamics and have emerged as a central platform in quantum thermodynamics and quantum technologies. Noting by the limitations of classical chemical batteries and the rapid miniaturization of quantum devices, it is natural to look towards QBs, which exploit genuinely quantum features such as quantum coherence, entanglement, and collective many-body dynamics to enhance performance beyond classical bounds [alicki2013entanglement, campaioli2017enhancing, seah2021quantum, Zhu2023Charging, downing2024hyperbolic, friis2018precision, Lu2025topological, rossini2020quantum, gyhm2022quantum, gyhm2024beneficial, rosa2020ultra, rodriguez2021collective, konar2024quantum, mazzoncini2023optimal, zhang2019powerful, yang2023battery, ferraro2018high, andolina2019quantum]. A wide range of theoretical strategies has been proposed to optimize charging protocols and performance of QBs [ferraro2018high, andolina2019quantum, PhysRevE.100.032107, rossini2020quantum, julia2020bounds, le2018spin, ghosh2020enhancement, PhysRevA.104.032207, PhysRevA.105.022628, PhysRevA.104.L030402, PhysRevB.102.245407, PhysRevResearch.4.033216, PhysRevA.106.022618, PhysRevA.110.022226, konar2024quantum, chaki2024positive, PhysRevA.106.062609, PhysRevA.107.032203, cyrc-ms34, chaki2024universal, PhysRevA.108.042618, PhysRevA.110.012227, PhysRevApplied.19.064069, PhysRevLett.132.090401, PhysRevLett.133.180401, mitra2025bound, perciavalle2025extractable, PhysRevLett.131.240401, PhysRevB.100.115142, PhysRevE.108.064106, gyhm2024beneficial, romero2025kicked, sahoo2025power, bukov2015universal, eckardt2017colloquium, Puri2024Floquet, Lu2025topological, ferraro2018high, andolina2019extractable, GarciaPintos2020], many of which have been realized experimentally in platforms such as quantum dots [PhysRevLett.131.260401], superconducting transmons [PhysRevA.107.023725, hu2022optimal, gemme2022ibm], organic semiconductors [quach2022superabsorption], superconducting circuits [hu2026quantum], and nuclear magnetic resonance systems [PhysRevA.106.042601]. In addition, QBs have been extensively studied in open quantum system settings, where environmental effects such as dissipation, decoherence, and engineered reservoirs strongly influence charging efficiency, stability, and extractable work [Quach2020, Tabesh2020EnvMediated, Morrone2022NonMarkov, Yao2025Reservoir, QuantumErgotropy2025, NonMarkovNSpin2025, Xu2023CompositeEnv, Zhang2024OpenInteractive, Zhang2024Open, EntropicUncertaintyQB2022, liu2025open].
Interacting many-body systems provide a natural platform for quantum batteries, where collective dynamics and correlations can significantly enhance charging power, accelerate energy storage, and improve work extraction compared to independent cells [binder2015quantacell, campaioli2017enhancing, ferraro2018high, andolina2019quantum, andolina2019extractable, le2018spin, rodriguez2021collective]. Such systems, including interacting spin chains, not only exhibit nontrivial scaling with system size but also serve as versatile platforms for probing nonequilibrium quantum thermodynamics and many-body dynamics [alicki2013entanglement, Puri2024Floquet, romero2025kicked]. Quantum battery architectures consist broadly of interacting batteries driven by noninteracting systems or noninteracting batteries charged by interacting many-body chargers, with performance strongly dependent on the Hamiltonian structure, interaction range, and relative strength of the charger and the battery [le2018spin, ghosh2020enhancement, zhao2021quantum, rossini2020quantum, ferraro2018high, julia2020bounds]. Motivated by this strong sensitivity to microscopic details, periodically driven and time-dependent protocols have attracted growing attention, as they enable the engineering of effective interactions, enhance robustness, and further improve charging performance [mondal2022periodically, romero2025kicked, bai2020floquet, Puri2024Floquet, downing2023quantum, crescente2020charging, shukla2026many]. In most existing studies, the charger and the battery are treated as two distinct subsystems, with energy transferred between them via an interaction. In contrast, in our setup we consider a single many-body system that plays the role of both the battery and the charger, governed by different Hamiltonians in each stage. Specifically, the system evolves under one Hamiltonian when acting as the battery and under another during the charging process. To distinguish these roles, we refer to the Hamiltonian governing the charging dynamics as the “charging Hamiltonian” (or “chargoid”), while the other defines the battery Hamiltonian.
Entanglement has long been considered a potential resource for enhancing quantum battery performance [ferraro2018high, PhysRevResearch.2.023113], as global collective driving protocols can generate entanglement among battery units and potentially increase instantaneous charging power compared to local parallel schemes [binder2015quantacell, campaioli2017enhancing, andolina2019quantum]. This led to the view that multipartite entanglement could be a key ingredient for fast charging. However, subsequent research has clarified that entanglement alone does not guarantee improved charging or work extraction; instead, performance is primarily determined by the structure of interactions and collective effects, rather than the presence of entanglement itself [PhysRevE.102.052109, le2018spin, gyhm2024beneficial].
Intrigued by this unresolved interplay between charging power and the buildup of correlations, we investigate the temporal relationship between energy flow and quantum correlations. To this end, we compare the instantaneous charging power with several entanglement-based measures (bipartite, tripartite, and multipartite correlations) by analyzing two-body nearest-neighbor interactions in both configurations (interacting battery with a non-interacting chargoid, and non-interacting battery with an interacting chargoid). Our results reveal a consistent temporal ordering, in which the instantaneous power attains its maximum before any of the considered entanglement measure quantities. This behavior indicates that entanglement-related correlations do not develop faster than energy storage, suggesting that entanglement alone does not play a decisive role in determining the stored energy of the system.
To elucidate the mechanism underlying this temporal separation, we analyze charging protocols governed by $\kappa$-local interactions under norm-constrained (“fair”) conditions. In $\kappa$-local schemes with $\kappa<N$, with NN being the system-size, entanglement is generated only within a restricted subset of particles, leaving the remaining degrees of freedom effectively inactive during the charging process; consequently, the resulting correlations cannot significantly contribute to a global charging advantage. In contrast, for fully collective interactions with $\kappa=N$, all particles are simultaneously involved in the entanglement dynamics, allowing correlations to directly support energy transfer. In this all-to-all interacting regime, the growth of entanglement and the instantaneous power exhibit similar temporal behavior, highlighting the essential role of genuinely global correlations in collective charging enhancement.
@@
The paper is organized as follows. Section II introduces the quantum battery and chargoid and defines measures for bipartite, tripartite, and multipartite entanglement. Section III presents the charging dynamics for both configurations: interacting battery with noninteracting chargoid, and noninteracting battery with interacting chargoid and examines the effects of $\kappa$-local interactions on energy transfer and correlations of particles. Section IV summarizes the main findings and concludes the study.
Figure 1: Schematic illustration of the two charging protocols. (a) The battery is initially non-interacting and charging is achieved by switching on spin–spin interactions at $t>0$, which act as the chargoid. (b) The battery consists of interacting spins and charging is implemented via independent local magnetic fields applied to each spin at $t>0$.
@@
Quantum Battery and chargoid
@@
To initiate the charging process, we switch on the interactions among the battery spins at time $t=0$, illustrated in Fig. 1(a). Additionally, we include an external field applied in the $x$-direction, which further drives the charging dynamics. The resulting evolution is governed by the interacting spin Hamiltonian
@@
where $J_{\alpha}$ are the nearest-neighbor interaction strengths in the $\alpha(\equiv x,y,z)$-direction, and $h_{x}$ is the magnetic field applied to each spin in the $x$-direction. $\kappa$ quantifies the number of interacting spins, representing the degree of $\kappa$-locality. By varying $\kappa$, we examine how higher-order spin correlations influence charging performance and the generation of multipartite entanglement. This Hamiltonian simultaneously induces spin correlations and drives the energy injection that charges the battery. The charging dynamics of the battery are governed by the Hamiltonian
@@
We also examine the behavior of the stored energy when the battery and chargoids are interchanged, as illustrated in Fig. 1(b).
@@
Concurrence is one of the most widely used measures of bipartite entanglement for two-qubit systems. It provides a quantitative value ranging from 0 for separable (uncorrelated) states to $1$ for maximally entangled states. In the context of quantum batteries, concurrence is useful for quantifying the entanglement between specific pairs of spins, for example, the first and last spins of a chain, which can reveal the formation of long-range correlations during the charging process.
For a two-qubit density matrix $\rho$, we define the spin-flipped matrix as [wootters1998entanglement, subrahmanyam2004quantum, osborne2002entanglement, osterloh2002scaling, arnesen2001natural] $\tilde{\rho}=(\sigma_{y}\otimes\sigma_{y})\rho^{*}(\sigma_{y}\otimes\sigma_{y})$, where $\rho^{*}$ is the complex conjugate in the computational basis and $\sigma_{y}$ is the Pauli-$y$ matrix. Using this, we construct $R=\rho\tilde{\rho}$. Let $\{\mu_{i}\}_{i=1}^{4}$ be the eigenvalues of $R$ (which are always real and nonnegative), and let their square roots be arranged in decreasing order: $\lambda_{1}\geq\lambda_{2}\geq\lambda_{3}\geq\lambda_{4}\geq 0$. The concurrence is then given by
$C(\rho)=\max\Big\{0,\lambda_{1}-\lambda_{2}-\lambda_{3}-\lambda_{4}\Big\}.$
@@
Bipartite entanglement entropy (BEE) quantifies the degree of entanglement between a subsystem $X$ and its complement $X^{c}$ in a many-body system. It is computed from the entropy of the reduced density matrix associated with the subsystem [hamma2005bipartite, PhysRevLett.96.181602, calabrese2004entanglement, page1993average, jiang2025entanglement, kadar2010entanglement],
@@
To study correlations among three subsystems simultaneously, we use the tripartite mutual information (TMI), defined as [hosur2016chaos, ding2016conditional, iyoda2018scrambling, seshadri2018tripartite, shukla2025scrambling, PhysRevB.100.224302]
@@
and its magnitude directly reflects the number of spins that are genuinely entangled. Specifically, for an $N$-spin battery, exceeding the threshold
$F_{Q}[\hat{H}_{B}]>\Big\lfloor\frac{N}{\kappa}\Big\rfloor\kappa^{2}+\Big(N-\Big\lfloor\frac{N}{\kappa}\Big\rfloor\kappa\Big)^{2}$ (9)
guarantees at least $(\kappa+1)$-partite entanglement, providing a rigorous lower bound on the depth of correlations [Naik2019Controlled].
Crucially, multipartite entanglement is a key resource in quantum batteries: it enables collective correlations among spins, allowing the system to store and transfer energy more efficiently than uncorrelated or weakly correlated states. A large QFI thus not only signals the presence of entanglement but also quantifies the extent to which the spins act coherently, offering insight into the fundamental limits of energy storage and extraction in many-body quantum systems. In this way, QFI serves as both a diagnostic tool for genuine quantum correlations and a benchmark for the performance potential of quantum batteries.
@@
While BEE measures correlations for a single partition, the average bipartite entanglement entropy (ABEE) provides a global view of entanglement across the system [igloi2008finite]. It is computed by averaging the BEE over all contiguous bipartitions:
@@
Figure 2: Noninteracting battery charged by an interaction-induced chargoid. Bipartite entanglement $(a_{1},b_{1},c_{1})$: Time evolution of the instantaneous power $P_{i}$, concurrence between the first and last spins $C(1,N)$, and BEE $S_{N/2}$. Tripartite entanglement $(a_{2},b_{2},c_{2})$: Time evolution of $P_{i}$ and TMI $I_{3}$. Multipartite entanglement $(a_{3},b_{3},c_{3})$: Time evolution of $P_{i}$, QFI $F_{Q}$, and ABEE $\overline{S}_{N/2}$. Panels correspond to: $(a_{1},a_{2},a_{3})$ $J_{x}=1$, $J_{y}=J_{z}=0$; $(b_{1},b_{2},b_{3})$ $J_{x}=J_{y}=1$, $J_{z}=0$; $(c_{1},c_{2},c_{3})$ $J_{x}=J_{y}=J_{z}=1$. Number of spins: $N=8$ and $h_{x}=h_{z}=1$.
@@ III Results
Noninteracting Battery with an Interaction-Induced chargoid
We consider a configuration in which the battery is initially a noninteracting spin system, described by the Hamiltonian $H_{B}$. Charging is implemented by turning on spin–spin interactions among the same spins, described by the Hamiltonian $H_{C}$, which act as the chargoid. These interactions are active only during the charging interval $0,\tau]$, driving energy flow into the battery while simultaneously generating quantum correlations (a similar setup was discussed in Ref. [alicki2013entanglement]). In this setup, the noninteracting battery serves as a passive energy receiver, allowing us to isolate and study the effects of chargoid-induced correlations on the charging dynamics.
@@
We investigate the relationship between the battery charging rate and bipartite correlations. To quantify bipartite correlations, we evaluate two complementary measures: the concurrence and the BEE. These quantities are computed for three representative chargoid interaction configurations: ($a_{1}$) $J_{x}=1$, $J_{y}=0$, $J_{z}=0$; ($b_{1}$) $J_{x}=1$, $J_{y}=1$, $J_{z}=0$; and ($c_{1}$) $J_{x}=1$, $J_{y}=1$, $J_{z}=1$. In all cases, the field strength is fixed to $h_{x}=1$. Concurrence quantifies pairwise entanglement between two qubits; here, we focus on the first and last spins to probe long-range correlations. In contrast, the BEE measures the entanglement between a chosen subsystem and its complement through the BEE of the reduced density matrix. In our calculations, we consider equal bipartitions of the spin chain. For a meaningful comparison, all three quantities are normalized by their respective maximum values, allowing us to identify the time regimes in which each quantity attains its peak.
@@
During the initial charging stage the instantaneous power $P_{i}(t)$ rises rapidly as the chargoid injects excitations into the battery; this rise is driven by coherent excitation transfer enabled by the chargoid’s internal couplings. Concurrence and BEE also increase during this phase, indicating that entanglement participates in the energy transfer process. Crucially, however, the peak of instantaneous power is reached before concurrence and BEE attain their maxima (see Fig. 2 $(a_{1},b_{1},c_{1})$). Concurrence typically grows faster than BEE and peaks earlier than BEE, but both lag the power peak. Physically, this temporal ordering shows that short-time coherent transport, which governs peak power, precedes the formation of fully developed bipartite entanglement across the system. Entanglement is therefore a necessary resource for enhanced charging, but it is not the sole determinant of the timing of maximal power: the chargoid’s coupling strengths and the coherent transfer channels set the time scale for the power peak, whereas entanglement continues to build up afterwards as correlations spread.
@@
To probe correlations beyond pairwise contributions, we evaluate the TMI, which quantifies how information and correlations are distributed among three subsystems. For this calculation, the spin chain of length $N$ is partitioned into four segments of size $N/4$ in total, and the initial state is taken to be the ground state of the battery Hamiltonian. This ensures consistency between the injected energy and the resulting correlation production during the charging process. The TMI increases substantially only after the early-time rise of energy and power, and its maximum occurs at later times compared to energetic quantities [Fig. 2 $(a_{2},b_{2},c_{2})$)]. This delayed growth of TMI reveals that three-way and higher-order correlations require additional time to form: pairwise coherences must first be established and propagated before genuine three-body information sharing emerges. From a physical standpoint, this hierarchical buildup reflects the propagation and interference of excitations through the interacting chargoid, leading ultimately to delocalized correlation structures (scrambling) that are slower to develop than local energy transfer.
@@
Energy injection from an interacting chargoid initially excites coherent transport channels that rapidly transfer energy into the battery and produce a pronounced peak in instantaneous power. The formation of quantum correlations pairwise, tripartite, and finally multipartite follows this energetic surge because correlations require the accumulation of relative phases and interaction-mediated spreading of excitations. Consequently, entanglement and scrambling develop on longer timescales than energy transfer. Therefore, although entanglement is essential to realize quantum advantages in charging (for instance by enabling nonclassical interference pathways or collective enhancement), there is no one-to-one temporal correspondence between entanglement maxima and the peak of instantaneous power. Instead, optimal charging occurs in an intermediate regime where coherence and modest entanglement combine to maximize power, while more extensive entanglement emerges later and contributes to the global reorganization and stabilization of the stored energy.
Interacting Battery with Noninteracting External-Field chargoid
We now consider a configuration in which the battery consists of interacting spins, while charging is implemented via an external field applied independently to each spin. Since the charging field does not induce interactions among the spins, it effectively acts as a noninteracting chargoid (similar scenarios are considered in Refs. [le2018spin, ghosh2020enhancement]). In this setup, the roles of the Hamiltonians are effectively reversed: the battery Hamiltonian [Eq. (1)] serves as the chargoid Hamiltonian $\hat{H}_{C}$, while the chargoid Hamiltonian [Eq. (2)] functions as the battery Hamiltonian $\hat{H}_{B}$. This setup allows us to investigate how the internal interactions within the battery influence energy transfer and dynamics, while isolating the effect of the chargoid from correlation buildup.